OpenStudy (andrewkaiser333):

I bought a new guitar for $800. It depreciates in value by 8% a year. Write an equation to model the great value. What will be its value in 5 years?

OpenStudy (andrewkaiser333):

This is what I got so far. G=800(.08)^n-1

OpenStudy (andrewkaiser333):

and then G=800(.08)^5-1

OpenStudy (retireed):

So what do you get for an answer? And does it make sense?

OpenStudy (andrewkaiser333):

the thing is i am not sure if i did this right

OpenStudy (retireed):

Well you need to test your equation to see if it seems logical. And thanks for sowing your thought process.

OpenStudy (retireed):

showing*

OpenStudy (andrewkaiser333):

np. right now this is what i remember doing a few day back for it but forgot whether i got it all of whether i am missing something.

OpenStudy (retireed):

So using your thought process, what would the the value of the guitar at the end of ONE year.

OpenStudy (andrewkaiser333):

4.12=G=800(.08)^2-1

OpenStudy (andrewkaiser333):

?

OpenStudy (retireed):

Your equation would be... G = 800(.08)^(1-1) which would be G= 800(0.08)^(1-1) = 800(0.08)^0 = 800(1) = 800 so after one year your equation would say NO depreciation, which can't be correct.

OpenStudy (andrewkaiser333):

oops sorry

OpenStudy (andrewkaiser333):

i gave year 2 by accident

OpenStudy (retireed):

The exponent ideal is correct, but there can't be any minus 1. So, just use 1 for one year or 5 for five years. Okay?

OpenStudy (andrewkaiser333):

ok so G=800(.08)^5-1=64^4=$16,777,216

OpenStudy (andrewkaiser333):

is that correct

OpenStudy (mathmale):

This is an alternative approach: If the guitar depreciates by 8% per year, then its value after 1 year would be (1.00-0.08)*800. Equivalently, value(1)=$800(0.92)^1 The value after 2 years would be value(2)=$800(0.92)^2. What would the value be after 5 years? value(5)=?

OpenStudy (retireed):

Do the exponent before the multiply by 800.

OpenStudy (mathmale):

Yes; that follows "order of operations" rules.

OpenStudy (retireed):

The equation format should look like ... G=800(some percentage depreciation)^N

OpenStudy (andrewkaiser333):

yes so did i do it right?

OpenStudy (retireed):

No

OpenStudy (andrewkaiser333):

i did the formula

OpenStudy (andrewkaiser333):

G=800(.08)^5-1

OpenStudy (retireed):

Do the exponent calculation before the multiply by 800. Remember PEDMAS.

OpenStudy (retireed):

G=800(some percentage depreciation)^N

OpenStudy (andrewkaiser333):

then 800(.08)=64^5-1

OpenStudy (mathmale):

Hold the $800!! Focus on (0.92)^5 first.

OpenStudy (andrewkaiser333):

oh

OpenStudy (mathmale):

You are attempting to evaluate this 800(0.08)^5 from left to right. DON'T. You MUST do the exponentiation first. What is (0.92)^5?

OpenStudy (andrewkaiser333):

(.08)^4=4.096e^-5

OpenStudy (andrewkaiser333):

4.096e^-5(800)

OpenStudy (mathmale):

resulting in ... ? Please note: the value at the beginning is $800. The value after 1 year is $800(0.92)^1. tHE VALUE after 5 years is $800(0.92)^5 (not 4).

OpenStudy (andrewkaiser333):

this =.032768

OpenStudy (mathmale):

Once more, Andrew: Please evaluate (0.92)^5

OpenStudy (andrewkaiser333):

yes but it is ^5-1

OpenStudy (mathmale):

You get 0.32768; I get 0.659.

OpenStudy (andrewkaiser333):

so that would be ^4

OpenStudy (andrewkaiser333):

this is why i am confused

OpenStudy (mathmale):

why? Andrew, please go back and read what I've typed. ^4 will give you the value of the guitar after 4 years. Is that what you want? Are you reading what I'm typing?

OpenStudy (andrewkaiser333):

i am trying tounderstand show me the work you did please so i see what you me

OpenStudy (mathmale):

Does this make sense to you? The value of the guitar at the beginning is $800(0.92)^0. What is (0.92)^0?

OpenStudy (andrewkaiser333):

1

OpenStudy (mathmale):

Everything I did I have typed in for your benefit. You'll need to re-read some of our conversation. Or, just answer my qeustion, above. Right. So, after 0 years, the value is $800(0.92)^0 = $800(1) = $800. Right or not?

OpenStudy (andrewkaiser333):

yes

OpenStudy (mathmale):

In other weords, the value right now is $800. Fine. What is the value after 1 year? $800(0.92)^1 = ??

OpenStudy (andrewkaiser333):

736

OpenStudy (mathmale):

Great. What is the value after 3 years?

OpenStudy (andrewkaiser333):

622.9504

OpenStudy (mathmale):

Right you are. What is the value afer 5 years?

OpenStudy (andrewkaiser333):

527.2652186

OpenStudy (mathmale):

Excellent. You're done. You're correct. Congrats. You do need to round that off to $527.27.

OpenStudy (andrewkaiser333):

ok

OpenStudy (mathmale):

Happy?

OpenStudy (andrewkaiser333):

so what would be the full equation for it?

OpenStudy (andrewkaiser333):

just 800(.08)^5?

OpenStudy (mathmale):

You can deduce that from the pattern we've used. The value of the guitar after 4 years is $800(0.92)^4. The value after x years is $800(0.92)^ ??

OpenStudy (andrewkaiser333):

even though the formula is 800(.08)^n-1

OpenStudy (mathmale):

That formula would be correct only if you use a different interpretation. Example: The value of the guitar at the beginning of year 1 is $800(0.92)^(1-1). What is this value?

OpenStudy (andrewkaiser333):

that is what the book showed me i was just trying to use this this is how i got confused i guess

OpenStudy (mathmale):

that's why it's so important to DEFINE YOUR TERMS.

OpenStudy (mathmale):

Book uses "at the beginning of year 1": $800(0.92)^(1-1) I use "at the end of the first year:" $800(0.92)^1

OpenStudy (andrewkaiser333):

so for the 5 years in this formula would be 800(.08)^(6-1)?

OpenStudy (mathmale):

In truth I should have followed your book's explanation, but I do not have a copy. To answer your question, you must define your formula. Are you talking about "at the beginning of year 1" or "at the end of year one"?

OpenStudy (andrewkaiser333):

?

OpenStudy (mathmale):

If you preface your previous statement with: "At the beginning of year 6, the value is..." and your formula will give you $800(0.92)^5 (which is corr4ect). If you preface

OpenStudy (mathmale):

if you preface your prevsious statement with "At the end of year 5, the value is $800(0.92)^5, you'll get the same (correct) result.

OpenStudy (andrewkaiser333):

so i am right?

OpenStudy (andrewkaiser333):

for that type of formula

OpenStudy (andrewkaiser333):

?

OpenStudy (mathmale):

Andrew, please read what I said. You have to explain your terms. Either you are talking about "the beginning of year x)" or "the end of year 5". Take your pick. The formulas are almost, but not quite, the same. Again I'm sorry I didn't have your book available, so I would have followed the book example instead of my own. But both formulas are correct IF AND ONLY IF you define your terms as explained above.

OpenStudy (andrewkaiser333):

can you try to use the formula i used from the book G=800(.08)^n-1

OpenStudy (andrewkaiser333):

if i do not use it i am going to get it wrong

OpenStudy (mathmale):

Yes. First, I define my perspective. "At the BEGINNING of year 1, the value of the guitar is \[$800(0.92)^{1-1}.\]

OpenStudy (andrewkaiser333):

ok

OpenStudy (mathmale):

At the beginning of year 3, the value is \[$800(0.92)^{3-1}, or. $800(0.92)^2\]

OpenStudy (andrewkaiser333):

yes

OpenStudy (andrewkaiser333):

so for year 5

OpenStudy (mathmale):

So after 5 years, we are at the beginning of year 6, right? would you please write the correct expression for the value then?

OpenStudy (andrewkaiser333):

G=800(.08)^(6-1)

OpenStudy (andrewkaiser333):

?

OpenStudy (mathmale):

Your formula uses the "at the beginning of year #" definition. Now imagine, we're at the end of year 5, which is the beginning of year 6. So, you MUST use 6 for n. What is the value of the guitar after 5 years?

OpenStudy (mathmale):

Regarding your "G=800(.08)^(6-1)" : What happened to our 0.92?

OpenStudy (andrewkaiser333):

it has been .08 the whole time

OpenStudy (andrewkaiser333):

8% not 92%

OpenStudy (mathmale):

Not so. Please go back and review our discussion. The guitar depreciates at the rate of 8% per year. True!! But if you want to know the value of the guitar in the beginning of the nth year, you write $800(1-0.08)^(n-1), or $800(0.92)^(n-1).

OpenStudy (andrewkaiser333):

?

OpenStudy (mathmale):

Depreciates = loses value Depreciation rate = 0.08 per year Factor with which to find Value AFTER depreciation = (1-0.08) = 0.92

OpenStudy (mathmale):

You calculated the value after 1 year. What was it? go back and find this value.

OpenStudy (andrewkaiser333):

idk i am too confused

OpenStudy (mathmale):

You found that it was $736 and you were right. Now I'd like for y ou to calculate $800(0.08)^1. Next, I'd like for you to calculate $800(1-0.08)^1, or $800(0.92)^1. please go ahead. Which one produces the correct answer, $736?

OpenStudy (andrewkaiser333):

the bottom two equations

OpenStudy (mathmale):

OK. One more thing. Calculate $800(0.08)^1 again.

OpenStudy (andrewkaiser333):

64

OpenStudy (mathmale):

Yes. That, Andrew, is the DEPRECIATION. wHAT IS $800-$64?

OpenStudy (andrewkaiser333):

800-64=736

OpenStudy (mathmale):

That's right. That's exactly what you got earlier. The avlue after 1 year is $736. That's found by taking the initial value, $800, and subtracting the depreciation, $64. Does this make things any clearer? Neither $800(0.08)^n nor $800(0.08)^(n-1) will give y ou the value of your guitar. But either $800(0.92)^(whatever) will.

OpenStudy (mathmale):

Where did that 0.92 come from? Answer: we subtract the annual depreciation rate (8%) from 100%. or subtract 0.08 from 1.00. Result: 0.92.